# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v1

@@ 4.3 Welfare analysis
We follow Krusell et al. (2010) in measuring the welfare change for black and white workers by calculating the change in average consumption equivalence after removing discriminatory hiring. [^9] Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$. Under an alternate economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state. We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$. With $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$. We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change. We aggregate over the counterfactual distribution rather than benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms. Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift. We present the average welfare change between black and white workers by productivity types and by wealth quintiles in Table 4.
Overall, the economy experiences a 5.8% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race. Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%. This is because of the rise in their job opportunities and higher wage rates. On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%. This is because of the decrease in their wage rate and the increase in the unemployment rate.
Mid-panels in Table 4 present the heterogeneous welfare change for black and white workers by productivity types. Though the welfare reduction of eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms. In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race. In comparison, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
@@
The bottom panels of Table 4 present the welfare change by wealth quintiles for black and white workers. Removing discriminatory firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race. The most considerable reduction in welfare (9.43%) happens for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
@@ 5 Business cycle dynamics
In this section, we first document the racial disparities over the business cycle. Then, we introduce aggregate uncertainty in our benchmark model to quantify the role of hiring discrimination in generating disparate dynamics in the business cycle. Lastly, we discuss the welfare implications of hiring discrimination over the business cycle.
@@ 5.1 Racial disparities over the business cycles
Table 5 summarizes the business cycle disparities in the labor market between black and white workers. The top panel presents the dynamics of the racial unemployment rate gap, and the bottom panel presents the dynamics of the average wage. Cyclicality is measured by the correlation between the variable and real GDP. Volatility is calculated by the standard deviations of the variable relative to the standard deviation of real GDP.
Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference. We then log the values and apply the HP filter with a smoothing parameter 1600 to extract cyclical components. We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions. It has a relative volatility of 9.71.
@@
Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage. This is because of the composition change of workers over the business cycles. Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994). For the same reason, we report the cyclical properties of the average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index. The wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics. The wage volatility is calculated by the relative standard deviations of the year-to-year change in wage from the wage regression. [^10] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors. Specifically, we follow Bils (1985), Solon et al. (1994), and Devereux (2001) and run the regression: $\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it}$ (22) where $Y_{t}$ refers to real GDP. Demographic variables $x_{it}$ include gender, education, labor market experience, and marital status. We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$. We measure cyclicality using coefficient $\beta$ from the regression. The volatility is the standard deviation of the year fixed effects, $\nu_{t}$. We then adjust the estimation to quarterly frequency. We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics. Black workers’ average wage is more volatile (2.54) than white workers (2.08).
@@ 5.2 Augmented model with aggregate shocks
We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2. The TFP shocks follows an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$. The output of the matched firm-worker pairs depends on the realizations of both aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
@@
We follow Cooley et al. (1995) and Boppart et al. (2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015. All other parameters stay the same as in Table 1. This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
We obtain the stochastic equilibrium of our model by using the sequence space method of Boppart et al. (2018). We first solve non-linearly for the perfect foresight transitions to a single small MIT shock, i.e., an unexpected shock to the aggregate TFP. We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock. Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycle as a linear combination of the impulse response and the TFP realizations. [^11] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system. In Appendix C, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
@@
Table 6 compares data and model moments on unemployment and wages. Similar to standard search models, the stochastic dynamics of our model are under the influence of the Shimer Puzzle (Shimer, 2005). The augmented benchmark model generates similar cyclicality in the unemployment rate gap as data but much less volatility in the unemployment rate gap. Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility. However, qualitatively, we have a higher wage cyclicality and volatility for black workers than for white workers.
@@ 5.3 Impact of hiring discrimination over business cycle
We compare the benchmark dynamics with the counterfactual model without p firms to examine the role of discriminatory hiring over aggregate shocks. [^12] As with Table 3, the counterfactual economy is a one-sector economy with only $np$ firms. We do not further adjust race-specific destruction shocks.
@@ 5.3.1 Impulse responses
Figure 3 plots the response of black and white workers’ labor market, wealth, and consumption outcomes after a 1% unanticipated expansionary TFP shock. We compare the benchmark model response to the counterfactual model without $p$ firms. Overall, our models generate countercyclical unemployment rates and procyclical wage, wealth, and consumption movements.
@@
In the model without $p$ firms, black workers’ unemployment rate decreases less than that of the benchmark economy after the expansionary shock. This implies that their unemployment rate is less volatile without $p$ firms. This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities to all workers. For white workers, the difference is much smaller between the benchmark and the model without $p$ firms. Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model. This is because the $p$ sector has a low job destruction shock. Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without discriminatory firms. This suggests more volatile procyclical wage and wealth responses for black workers than for white workers with the presence of $p$ firms. Lastly, We don’t observe a discerning difference in the consumption responses with and without $p$ firms.
@@ 5.3.2 Stochastic simulation
The counterfactual equilibrium without $p$ firms eliminates the racial unemployment rate gap, resulting in the disappearance of its cyclical properties. Similarly, the racial differences in the cyclicality and volatility of average wages are largely reduced. As shown in Table 7, in the benchmark model, the black-white unemployment rate gap exhibits strong countercyclicality (-0.6731) and volatility (0.3009), both of which vanish in the absence of $p$ firms. For average wages, the cyclicality for black workers decreases from 1.0749 in the benchmark to 1.0131 without $p$ firms, while for white workers it increases slightly from 0.9984 to 1.0115. Similarly, the volatility for black workers falls from 1.0755 to 1.0139, while for white workers it increases from 0.9985 to 1.0116. This convergence demonstrates that hiring discrimination is a primary driver of the labor market’s cyclical disparities between black and white workers. These findings align with Cajner et al. (2017), who argue that demographic differences alone cannot fully account for observed racial disparities over the business cycle.
@@
Note: The top panel compares the black-white unemployment rate, wage, consumption, and wealth gaps over the business cycles between the benchmark model and model without $p$ firms.
We further examine the behaviors in the cyclical properties of the average black-to-white wealth ratio and the average black-to-white consumption ratio. Other than the cyclicality of the average wealth ratio, the cyclical properties of the racial gaps in all other dimensions are reduced after discriminatory hiring is removed.
@@ 5.4 Welfare disparities in business cycles
Note: This table compares the average consumption equivalence change from the benchmark model to the one without $p$ firms in business cycles. We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type. Low, mid, and high productivity corresponds to the lowest, middle, and highest value of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The economy experiences an average of 0.01% welfare decrease after removing prejudiced firms. This is much less than the welfare change in the steady state in Table 4, primarily because workers are risk averse. All workers experience a decrease in unemployment rate volatility, and white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions. On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%. Among white workers, those with high productivity and those in the top 20% of the wealth distribution lose the most (0.02%). In contrast, black workers in the top wealth quintile benefit the most, with an average welfare gain of 0.03%.
@@ 6 Conclusion
This study examines the effects of racial hiring discrimination within a frictional labor market on employment, wage disparities, and wealth accumulation between black and white workers in the U.S. We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation. Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in steady state and business cycles. Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum. Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome within frictional markets. Eliminating these discriminatory factors results in a decrease in overall welfare. Our findings shed light on the enduring nature of Black-White racial disparities in the U.S., offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
@@ Appendix A Stationary Equilibrium
    Unemployment insurance tax rate $\tau$ solves to balance the government budget.
@@ Appendix B Comparing the impact of bargaining power and extreme wealth shock
Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks. Equalizing black workers’ bargaining power to white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes. This is because non-prejudiced firms retain less profit, hence posting fewer vacancies. On the net, it also reduces white worker’s job-finding rate and average wage rate. As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
@@
In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers. The effect resembles assigning a higher bargaining power to black workers. This is because of the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012). With a lower probability of losing their wealth, black workers can accumulate more personal wealth. Higher personal wealth gives black workers higher reservation value when bargaining with firms. Effectively, black workers can bargain for higher wage outcomes (1.66 compared to 1.60 in the benchmark model). Similar to directly assigning a higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent. The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output. It spills over to an increase in the average wage rate for all. Though it has a much less impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap. Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).

### v2

@@ 4.3 Welfare analysis
We follow Krusell et al. (2010) in measuring the welfare change for black and white workers by calculating the change in average consumption equivalence after removing discriminatory hiring. [^11] Under the benchmark model, let $V(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(c_{t})$ be the maximal value of the individual with employment status $e$, race $R$, productivity $s$, and asset $a$. Under an alternate economy, let $\tilde{V}(e,R,s,a)=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$ be the maximal value of individuals with each corresponding state. We examine the welfare change between the two economies through consumption equivalence $\Omega$, following the equation: $\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log((1+\Omega)c_{t})=\mathrm{E}_{0}\sum_{t=0}^{\infty}\beta^{t}log(\tilde{c}_{t})$. With $\log$ utility, we derive $\Omega=\exp((\tilde{V}-V)(1-\beta))-1$. We aggregate the individual-level consumption equivalence, $\Omega$s, using the distribution of the counterfactual economy to calculate the average welfare change. We aggregate over the counterfactual distribution rather than the benchmark distribution because eliminating $p$ firms in the model removes the distribution of white workers associated with the states of working at and unemployment from $p$ firms. Aggregating over the benchmark distribution overstates the welfare change for white workers without accounting for the distribution shift. We present the average welfare changes for black and white workers by productivity types and wealth quintiles in Table 5.
Overall, the economy experiences a 5.83% reduction in average welfare if we eliminate $p$ firms and a 1.17% increase in average welfare if we separate the job destruction shocks by race. Once we remove discriminatory hiring, black workers experience an average welfare increase of between 6.89% and 10.94%. This is because of the rise in their job opportunities and higher wage rates. On the other hand, white workers encounter an average welfare loss of between 1.14% and 8.83%. This is because of the decrease in their wage rates and the increase in the unemployment rate.
Mid-panels in Table 5 present the heterogeneous welfare change for black and white workers by productivity types. Though the welfare reduction from eliminating prejudiced firms is similar across productivity for white workers, the most considerable loss is for the lowest productive white workers at 9.03% when we remove $p$ firms. In comparison, the highest productive workers lose the most at 1.19% when further adjusting the job destruction rate by race. Furthermore, the most productive black workers have the highest welfare gain at 7.41% and 11.49% across the two counterfactual models.
@@
The bottom panels of Table 5 present the welfare change by wealth quintiles for black and white workers. Removing prejudiced firms benefits the mid-quintile black workers the most, at 6.95%, and the lowest-quintile black workers the most, at 10.98%, after adjusting the job destruction shocks by race. The most considerable reduction in welfare (9.43%) occurs for the lowest quintile white workers in the model removing $p$ firms and for mid-quintile white workers at 1.17% after adjusting for race-specific job destruction shocks.
@@ 5 Business cycle dynamics
In this section, we begin by documenting the racial disparities over business cycles. Next, we incorporate aggregate uncertainty into our benchmark model to quantify the impact of hiring discrimination on generating racially disparate dynamics within these cycles. Finally, we examine the welfare consequences of hiring discrimination during business cycles.
@@ 5.1 Racial disparities over the business cycles
Table 6 presents the business cycle disparities in labor market outcomes between black and white workers. The top panel shows the dynamics of the racial unemployment rate gap, and the bottom panel shows the dynamics of the average wage. Cyclicality is measured by the correlation between the variable and real GDP. Volatility is calculated by the standard deviation of the variable relative to the standard deviation of real GDP.
Unemployment rate: We use monthly data from the Current Population Survey (CPS) from 1996 to 2014 to calculate the black-to-white unemployment rate difference. We then log the values and apply the HP filter with a smoothing parameter of 1600 to extract cyclical components. We find that the racial unemployment rate gap is countercyclical, widening during recessions and narrowing during expansions. It has a relative volatility of 9.71.
@@
Wage: We rely on the individual data in PSID from 1996 to 2014 to construct a wage panel to calculate the cyclical properties of the average wage. This is due to the composition change of workers over the business cycles. Using aggregate wage data underestimates its cyclical properties (Stockman, 1983; Bils, 1985; Solon et al., 1994). For the same reason, we report the cyclical properties of average wages for black and white workers separately rather than calculating the cyclical properties for a wage differential index. Wage cyclicality is calculated through a wage regression as the percentage change of average wage when real GDP increases by one percent after controlling for demographics. Wage volatility is calculated by the relative standard deviations of the year-to-year change in wages from the wage regression. [^12] We conduct wage regression by regressing the change in individual log wage on the change in log GDP, after controlling for demographic factors. Specifically, we follow Bils (1985), Solon et al. (1994), and Devereux (2001) and run the regression: $\Delta\log w_{it}=\alpha+\beta\Delta\log Y_{t}+\gamma x_{it}+\nu_{t}+\upsilon_{s}+\epsilon_{it}$ (22) where $Y_{t}$ refers to real GDP. Demographic variables $x_{it}$ include gender, education, labor market experience, and marital status. We also include year fixed effects, $\nu_{t}$, and state fixed effects, $\upsilon_{s}$. We measure cyclicality using coefficient $\beta$ from the regression. The volatility is the standard deviation of the year fixed effects, $\nu_{t}$. We then adjust the estimation to quarterly frequency. We find that wages for black workers are twice as cyclical as wages for white workers once we control for demographics. Black workers’ average wage is more volatile (2.54) than that of white workers (2.08).
@@ 5.2 Augmented model with aggregate shocks
We introduce an aggregate total factor productivity (TFP) shock $z$ to the benchmark steady-state model from Section 2. The TFP shocks follow an AR(1) process, $\log(z^{\prime})=\rho_{z}\log(z)+\epsilon_{z}$, with $\epsilon_{z}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{z}^{2})$. The output of the matched firm-worker pairs depends on the realizations of aggregate and idiosyncratic productivity shocks and is given by $zsf(k)$.
@@
We follow Cooley et al. (1995) and Boppart et al. (2018) and choose the persistence parameter $\rho_{z}$ to be 0.95, while the standard deviation $\sigma_{z}$ is set to 0.015. All other parameters remain the same as in Table 1. This gives us an aggregate output volatility of 0.02 and a first-order auto-correlation of 0.73.
We obtain the stochastic equilibrium of our model by using the sequence space method of Boppart et al. (2018). We first solve non-linearly for the perfect foresight transitions to a single small MIT shock, i.e., an unexpected shock to the aggregate TFP. We then use the solved impulse responses as the numerical derivatives with respect to the initial TFP shock. Using these derivatives, we simulate the stochastic equilibrium by generating the TFP realizations and calculating the corresponding model moments over the business cycles as a linear combination of the impulse response and the TFP realizations. [^13] This method hinges on the assumption that the business cycle dynamics can be well approximated as a linear system. In Appendix C, we demonstrate the validity of this assumption by establishing the symmetry of the impulse responses to 1% positive and negative TFP shocks.
@@
Table 7 compares data and model moments on unemployment and wages. Similar to standard search models, the stochastic dynamics of our model are subject to the Shimer Puzzle (Shimer, 2005). The augmented benchmark model generates a similar cyclicality in the unemployment rate gap as the data but with much less volatility in the unemployment rate gap. Regarding the average wage, the model has lower cyclicality for both black and white workers and less pronounced differences in volatility. However, qualitatively, we observe a higher wage cyclicality and volatility for black workers than for white workers.
@@ 5.3 Impact of hiring discrimination over business cycle
We compare the benchmark dynamics with the counterfactual model without $p$ firms to examine the role of discriminatory hiring in the context of aggregate fluctuations. [^14] As with Table 4, the counterfactual economy is a one-sector economy with only $np$ firms. We do not further adjust race-specific destruction shocks.
@@ 5.3.1 Impulse responses
Figure 3 illustrates the labor market, wealth, and consumption responses of black and white workers following a 1% unexpected expansionary TFP shock. We evaluate the benchmark model’s reaction against the counterfactual model that excludes $p$ firms. In general, our models produce countercyclical unemployment rates alongside procyclical patterns in wages, wealth, and consumption.
@@
In the model without $p$ firms, the unemployment rate of black workers decreases less than that of the benchmark economy after the expansionary shock. This implies that their unemployment rate is less volatile without $p$ firms. This is because $np$ firms have to compete against $p$ firms for hiring, so they post more vacancies, providing more opportunities for all workers. For white workers, the difference between the benchmark and the model without $p$ firms is much smaller. Interestingly, their recovery in the counterfactual model is much faster than in the benchmark model. [^15] This is a common property in the literature on dual labor markets. A single-market model often reverts to its steady state faster than a dual-market model (e.g. Horvath and Yang, 2022). This is because the $p$ firms have a low job destruction shock. Once hired, white workers are less likely to lose their jobs, hence a slower rate to a steady state.
The average black-white wage and wealth ratios increase more after the expansionary shock in the benchmark model than without prejudiced firms. This suggests more volatile procyclical wage and wealth responses for black workers than for white workers in the presence of $p$ firms. Lastly, we don’t observe a discerning difference in the consumption responses with and without $p$ firms.
@@ 5.3.2 Stochastic simulation
In a counterfactual equilibrium without $p$ firms, the racial unemployment rate gap closes, eliminating its cyclical characteristics. Likewise, the disparities in the cyclicality and volatility of average wages between racial groups are significantly diminished. According to Table 8, the benchmark model reveals a pronounced countercyclicality (-0.6731) and volatility (0.3009) for the black-white unemployment rate gap, both of which disappear when $p$ firms are not included. For average wages, the cyclicality for black workers lowers from 1.0749 in the benchmark to 1.0131 without $p$ firms, whereas it increases slightly for white workers from 0.9984 to 1.0115. Conversely, the volatility for black workers drops from 1.0755 to 1.0139, while for white workers, it rises from 0.9985 to 1.0116. This convergence indicates that hiring discrimination significantly contributes to the cyclical disparities in the labor market between black and white workers. These results are consistent with Cajner et al. (2017), Ganong et al. (2020), and Ragusett (2022), which argue that demographic differences alone do not sufficiently explain the racial disparities observed throughout the business cycle.
@@
Note: The top panel compares the black-white unemployment rates, wages, consumption, and wealth gaps over the business cycles between the benchmark model and the model without $p$ firms.
@@ 5.4 Welfare disparities in business cycles
Note: This table compares the average consumption equivalence change from the benchmark model to the one without $p$ firms in business cycles. We simulate the economy for 10,000 periods and calculate the long-run average welfare for each individual type. Low, mid, and high productivity correspond to the lowest, middle, and highest values of idiosyncratic productivity $s$. The wealth quintiles are based on benchmark steady-state wealth distribution.
The economy experiences an average welfare decrease of 0.01% after removing prejudiced firms. This is much less than the welfare change in the steady state in Table 5, primarily because workers are risk averse. All workers experience a decrease in unemployment rate volatility, while white workers experience an increase in wage volatility that somewhat counters the changes in other dimensions. On average, black workers experience a modest welfare gain of 0.02%, whereas white workers face a slight welfare loss of 0.01%. Among white workers, those with high productivity and those in the top 20% of the wealth distribution lose the most (0.02%). In contrast, black workers in the top wealth quintile benefit the most, with an average welfare gain of 0.03%.
@@ 6 Conclusion
This study examines the effects of racial hiring discrimination within a frictional labor market on employment, wage disparities, and wealth accumulation between black and white workers in the U.S. We develop a search-and-matching model incorporating firms with and without racial prejudices alongside race-specific pathways for wealth accumulation. Our findings reveal that racial discrimination in hiring significantly exacerbates wage, unemployment, and wealth gaps in the steady state and business cycles. Moreover, discriminatory hiring disproportionately consigns black workers to the lower end of the wealth spectrum. Contrary to conventional discrimination theories, our analysis suggests that discriminatory hiring persists as an equilibrium outcome within frictional markets. Eliminating these discriminatory factors results in a decrease in overall welfare. Our findings shed light on the enduring nature of Black-White racial disparities in the U.S., offering insights into the interplay between discrimination, labor market dynamics, and wealth accumulation.
@@ Appendix A Stationary Equilibrium
    Labor income tax rate $\tau$ solves to balance the government budget of the total unemployment insurance payout.
@@ Appendix B Comparing the impact of bargaining power and extreme wealth shock
Table B.1 compares the benchmark model to models that remove non-market disparities, particularly in equalizing racial bargaining power and extreme wealth shocks. Equalizing black workers’ bargaining power with that of white workers (Column 2, $\xi_{bl}=\xi_{wh}$) depresses black workers’ job-finding rate but directly raises their bargained wage outcomes. This is because non-prejudiced firms retain less profit, hence post fewer vacancies. On the net, it also reduces white workers’ job-finding rate and average wage rate. As a result, the average wage gap shrinks to 19% (with a wage ratio of 81%), explaining about 24% of the racial wage gap.
@@
In Column 3 ($\epsilon_{bl}=\epsilon_{wh}$), we assign black workers the same conditions in accumulating wealth as white workers. The effect resembles assigning higher bargaining power to black workers. This is due to the importance of personal wealth in partial self-insurance against uncertain adverse outcomes (Nakajima, 2012). With a lower probability of losing their wealth, black workers can accumulate more personal wealth. Higher personal wealth gives black workers a higher reservation value when bargaining with firms. Effectively, black workers can negotiate for better wage outcomes (1.66 compared to 1.60 in the benchmark model). Similar to directly assigning higher bargaining power, the resulting job-finding rates decrease, though to a smaller extent. The smaller extreme wealth destruction rate also translates into better aggregate capital accumulation and increased aggregate output. It spills over to an increase in the average wage rate for all. Though it has a much lesser impact on the unemployment rate gap, it raises the average black-to-white wage ratio to 77% and reduces the wage gap to 23%, explaining 8% of the racial wage gap. Since it directly raises the wealth position of black workers, the mean and median black-to-white wealth ratios increase drastically (77% and 88% from the benchmark 28% and 33%).